# PVPO: Pre-Estimated Value-Based Policy Optimization for Agentic Reasoning

### v1

@@ Abstract
Critic-free reinforcement learning methods, particularly group policies, have attracted considerable attention for their efficiency in complex tasks. However, these methods rely heavily on multiple sampling and comparisons within the policy to estimate advantage, which may cause the policy to fall into local optimum and increase computational cost. To address these issues, we propose PVPO, an efficient reinforcement learning method enhanced by an advantage reference anchor and data pre-sampling. Specifically, we use the reference model to rollout in advance and employ the calculated reward score as a reference anchor. Our approach effectively corrects the cumulative bias introduced by intra-group comparisons and significantly reduces reliance on the number of rollouts. Meanwhile, the reference model can assess sample difficulty during data pre-sampling, enabling effective selection of high-gain data to improve training efficiency. Experiments conducted on nine datasets across two domains demonstrate that PVPO achieves State-Of-The-Art (SOTA) performance. Our approach not only demonstrates robust generalization across multiple tasks, but also exhibits scalable performance across models of varying scales.
@@ 1 Introduction
  • We introduce a group sampling strategy that offline filters data with unstable accuracy rates to construct high-quality batches, thereby enhancing convergence and learning efficiency. Furthermore, for samples with zero accuracy, we leverage a large-scale LLM to generate ground-truth trajectories, facilitating more effective learning from sparse reward signals.
  • PVPO achieves state-of-the-art performance on multi-step retrieval datasets and demonstrates strong generalization on mathematical reasoning benchmarks. Experimental results indicate that PVPO not only enhances multi-hop question answering and tool-use capabilities, but also improves the overall reasoning ability of LLMs.
@@ 4.1 Static V Estimate
$\begin{split}\hat{A}_{\text{dyn}}(\tau_{i},s_{0})&=\hat{Q}_{\text{dyn}}(\tau_{i})-\hat{V}_{\text{dyn}}(s_{0})\\ &=r_{i}-\text{mean}(\mathbf{r})\text{.}\end{split}$ (9)
@@
This stable static baseline replaces the unstable dynamic baseline in formula 8. We finally obtain the advantage function of PVPO, which is applicable to all sparse reward fields:
$\begin{split}\hat{A}^{\text{PVPO}}(\tau_{i},s_{0})&=\hat{Q}_{\text{dyn}}(\tau_{i})-\hat{V}_{\text{sta}}(s_{0})\\ &=r_{i}-\text{mean}(\mathbf{r^{\text{ref}}})\text{.}\end{split}$ (11)
For any step $(s_{i,t},a_{i,t})$ of any trajectory $\tau_{i}$, the advantage can also be written as follows:
$\hat{A}^{\text{PVPO}}_{i,t}$ $=r_{i}-\text{mean}(\mathbf{r}^{\text{ref}})\text{,}$ (12)
$r_{i}$ $=\begin{cases}\max(0.1,\,r_{\text{acc}}),&\text{if format is correct.}\\ 0,&\text{if format is incorrect.}\end{cases}$ (13)
$r_{\text{acc}}$ $=\begin{cases}\text{F}_{1}(a_{\text{pred}},\,a_{\text{gt}}),&\text{if }L_{\text{pred}}\geq n\cdot L_{\text{gt}}.\\ \text{CEM}(a_{\text{pred}},\,a_{\text{gt}}),&\text{if }L_{\text{pred}}<n\cdot L_{\text{gt}}.\end{cases}$ (14)
where $a_{*}$ denotes the answer, $L_{*}$ denotes the length of the answer, $_{pred}$ denotes the prediction, and $_{gt}$ denotes the ground truth. $r^{ref}_{i}$ is similar. F1 denotes the standard word-level F1 score and CEM denotes Cover Exact Match. $n$ represents the multiple of the text length, which we set to $3$ by default. This reward is used for agentic search, following Hao et al. (2025).
@@ 5 Expriments Setting
5.1 Metrics
For multi-hop question-answering tasks, we employ answer accuracy (Acc, %) and LLM-as-a-Judge (LasJ, %) (Song et al., 2025) as benchmarks. For mathematical reasoning tasks, we employ answer accuracy (Acc, %) as the evaluation metric, specifically reporting the mean accuracy across 32 independent rollouts for each sample (mean acc@32).
5.2 Data Collection and Processing
For multi-hop question-answering tasks, we conducted extensive experiments on four multi-step retrieval datasets, including Musique (Trivedi et al., 2022), 2WikiMultiHopQA (2Wiki) (Ho et al., 2020), HotpotQA (Yang et al., 2018), and Bamboogle (Bam) (Press et al., 2023). We trained on Musique, with the original dataset size of 20k and a filtered size of 8k after Group Sampling. Evaluations were done on the full development or test sets of these datasets.
For mathematical reasoning tasks, we trained on DAPO-Math-17k-Processed (Yu et al., 2025), with an original size of 17k and a filtered size of 10k after Group Sampling. We evaluated on five test sets: DAPO-AIME-2024 (AI-MO, 2024; Bytedance & Tsinghua-SIA, 2025), AIME-2025 (Lin, 2025), MATH500 (Lightman et al., 2024; HuggingFaceH4, 2023), AMC23 (AI-MO, 2024), and Olympiad (He et al., 2024).
5.3 Environment and Hyperparameters
All experiments ran on a server with an Intel(R) Xeon(R) Platinum 8369B CPU and eight NVIDIA A100-SXM4-80GB GPUs. We used Qwen2.5-7B-Instruct as the baseline model and Qwen2.5-72B-Instruct as the large LLM to generate GT Traj. $R^{\text{ref}}$ was updated every 500 steps. The number of pre-samples $M$ in the reference model was set equal to the number of samples $N$ in the policy model.
Our multi-hop question answering task used the ReSearch (Chen et al., 2025) framework, with batch size 8, rollout count $N=5$, and 1,000 training steps. DynaSearcher added “kg_filter” during inference, but we used direct inference throughout.
The mathematical reasoning task used the veRL (Sheng et al., 2025) framework, with batch size 32, rollout count $N=16$, and 1,000 training steps. GRPO’s “loss_agg_mode” was set to the default “token-mean” instead of “seq-mean-token-mean”. For evaluation, we set “val_temperature” = 0.6 and “val_top_p” = 0.95.
@@ 6 Experiments
In this section, we run a series of experiments. First, we test our method on multi-hop question answering to validate its effectiveness in the agent domain. Then, we test it on mathematical reasoning to show its generalizability. These experiments help us answer the following questions:
Q1: Compared to existing RL methods, can PVPO achieve better performance?
Q2: Does PVPO show generalizability across different domains?
Q3: Can PVPO reduce the temporal resource consumption required for training?
Q4: Does PVPO have better stability than baseline methods during training?
Q5: Does training with Static V Estimate really help the model?
6.1 Main Performance Comparison
To answer Q1, we evaluated PVPO on multi-step retrieval datasets and compared the results with two types of benchmarks. The results are shown in Table 1.
Table 1: Performance comparisons between PVPO and the baselines on multi-step retrieval datasets. The best and second best results are bold and underlined, respectively. $\dagger$$\ddagger$} represents in-domain/out-of-domain datasets.
Musique 2Wiki HotpotQA Bamboogle Avg Acc
@@ 6.1 Main Performance Comparison
Qwen2.5-7B-Instruct 5.10 13.48 27.90 29.30 22.40 31.04 12.80 17.12 17.05 22.74
DeepSeek-R1 32.00 40.70 57.50 59.40 43.00 58.30 66.40 76.64 49.73 58.76
O4-mini 38.00 44.10 61.50 67.40 49.50 67.40 74.40 84.16 55.85 65.77
GPT-4.1-global 31.00 40.90 58.00 58.50 44.50 57.70 51.20 61.60 46.18 54.68
Gemini-2.5-pro 42.50 50.80 70.00 71.20 53.00 71.10 75.20 84.48 60.18 69.40
@@
ReSearch 13.40 24.90 41.70 43.26 33.40 47.26 36.00 43.20 31.13 39.66
GRPO on ReSearch 33.40 46.72 60.80 67.02 54.50 63.68 45.60 54.40 48.58 57.96
GRPO on DynaSearcher 38.90 52.04 74.30 76.77 62.70 68.32 51.20 58.72 56.78 63.96
PVPO on ReSearch 36.50 51.44 70.10 72.36 65.50 72.34 45.60 54.32 54.43 62.62
PVPO on DynaSearcher 46.90 59.44 77.70 80.62 69.00 78.44 50.40 59.68 61.00 69.55
6.1.1 Performance of LLMs
We compared PVPO with the zero-shot inference performance of leading open-source and closed-source LLMs, including GPT-4.1-0414-global, DeepSeek-R1-0528, O4-mini-0416-global, and Gemini-2.5-pro-0325. The results show that the 7B model trained with PVPO improved significantly, reaching 3.6 times the accuracy of the original model. It outperforms all general LLMs on in-domain datasets. On the out-of-domain dataset, its performance is not as high as in zero-shot settings. This may be due to differences in training data and model capacity, since these leading models might have seen similar data during pre-training or fine-tuning, while the 7B model has limited capacity. However, its average accuracy is 8 percentage points higher than the average of the four LLMs and about 1 percentage point higher than the best-performing LLM.
6.1.2 Performance of Agentic Retrieval Methods
To enable a more fair comparison, we also compared PVPO with agentic retrieval methods that require training. In this experiment, all methods were trained for 1,000 steps. We included the direct inference methods of ReSearch and DynaSearcher, as well as a baseline that trained these two methods using RL with GRPO. We observed that the performance of PVPO improved by more than 5 percentage points on average compared to GRPO. The result demonstrates the consistent superiority of PVPO across agentic retrieval methods.
Overall, this answers Q1: Compared to existing reinforcement learning methods, PVPO achieves state-of-the-art performance.
Table 2: Performance comparisons between PVPO and GRPO under different scale models on mathematical reasoning datasets.
MATH500 AMC23 Olympiad AIME-2024 AIME-2025 Avg Acc
Qwen2.5-7B-Instruct
GRPO 78.66 49.20 42.26 14.26 12.83 39.44
PVPO 80.30 52.02 44.62 14.86 14.70 41.30
Qwen2.5-14B-Instruct
GRPO 82.22 55.94 49.70 17.14 16.90 44.38
PVPO 83.64 56.78 50.72 19.24 17.74 45.62
@@ 6.2 Generalization Evaluation
We shifted from multi-hop question answering to single-shot mathematical reasoning. We chose several math reasoning datasets with varying difficulty, from basic arithmetic to challenging olympiad-level problems. The goal is to test general reasoning ability in different fields and situations. In the experiment, we compared PVPO with GRPO. The results are shown in Table 2. In terms of average performance, PVPO outperforms GRPO at both model scales. On the 7B model, PVPO scores 1.89 percentage points higher than GRPO. When scaled up to 14B, PVPO continues to lead. PVPO shows clear advantages on most datasets.
Overall, PVPO achieved the best results on all five math reasoning datasets. This answers Q2: PVPO is generalizable and keeps stable performance across different fields and tasks.
Figure 2: Training efficiency study of PVPO on mathematical reasoning datasets.
@@ 6.3 Training Efficiency Analysis
6.3.1 Convergence Speed
We compare PVPO and GRPO to show PVPO’s faster convergence speed. During training, we tracked accuracy and rewards, as shown in Figure 2. Figure 2 (a) shows that after about 500 steps, PVPO reaches the accuracy level that GRPO gets after 1,000 steps. Figure 2 (b) shows that PVPO starts with a higher reward than GRPO and grows at a faster rate. Group Sampling provides high-gain samples during training, and the static advantage baseline gives a clear direction for policy optimization. Overall, PVPO reaches the target accuracy with fewer steps and converges faster.

### v3

@@ Abstract
Critic-free reinforcement learning methods, particularly group policies, have attracted considerable attention for their efficiency in complex tasks. However, these methods rely heavily on multiple sampling and comparisons within the policy to estimate advantage, which may cause the policy to fall into local optimum and increase computational cost. To address these issues, we propose PVPO, an efficient reinforcement learning method enhanced by an advantage reference anchor and data pre-sampling. Specifically, we use the reference model to rollout in advance and employ the calculated reward score as a reference anchor. Our approach effectively corrects the cumulative bias introduced by intra-group comparisons and significantly reduces reliance on the number of rollouts during training. Meanwhile, the reference model can assess sample difficulty during data pre-sampling, enabling effective selection of high-gain data to improve training efficiency. Moreover, PVPO is orthogonal to other advanced critic-free RL algorithms, making it compatible with and complementary to these methods. Experiments conducted on nine datasets across two domains demonstrate that PVPO achieves State-Of-The-Art (SOTA) performance. Our approach not only demonstrates robust generalization across multiple tasks, but also exhibits scalable performance across models of varying scales.
@@ 1 Introduction
  • We introduce a group sampling strategy that offline filters data with unstable accuracy rates to construct high-quality batches, thereby enhancing convergence and learning efficiency. Furthermore, for samples with zero accuracy (i.e., zero reward), we leverage a large-scale LLM to generate ground-truth trajectories, facilitating more effective learning from sparse reward signals.
  • PVPO achieves state-of-the-art performance on multi-step retrieval datasets and demonstrates strong generalization on mathematical reasoning benchmarks. Experimental results indicate that PVPO not only enhances multi-hop question answering (QA) and tool-use capabilities, but also improves the overall reasoning ability of LLMs.
@@ 4.1 Static V Estimate
$\begin{split}\hat{A}_{\text{dyn}}(\tau_{i},s_{0})&=\hat{Q}_{\text{dyn}}(\tau_{i})-\hat{V}_{\text{dyn}}(s_{0})=r_{i}-\text{mean}(\mathbf{r})\text{.}\end{split}$ (9)
@@
This stable static baseline replaces the unstable dynamic baseline in formula 8. We finally obtain the advantage function of PVPO, which is well-suited for RL tasks with sparse rewards.
$\begin{split}\hat{A}^{\text{PVPO}}(\tau_{i},s_{0})&=\hat{Q}_{\text{dyn}}(\tau_{i})-\hat{V}_{\text{sta}}(s_{0})=r_{i}-\text{mean}(\mathbf{r^{\text{ref}}})\text{.}\end{split}$ (11)
@@ 5 Expriments Setting
Metrics. For multi-hop QA tasks, we employ answer accuracy (Acc, %) and LLM-as-a-Judge (LasJ, %) (Song et al., 2025) as evaluation metrics. For mathematical reasoning tasks, we measure answer accuracy (Acc, %), reporting the mean accuracy across 32 independent rollouts for each sample (i.e., acc@32).
Datasets. For multi-hop QA tasks, we conduct experiments on four multi-step retrieval datasets: Musique (Trivedi et al., 2022), 2WikiMultiHopQA (2Wiki) (Ho et al., 2020), HotpotQA (Yang et al., 2018), and Bamboogle (Bam) (Press et al., 2023). Model training is performed on the Musique training split, which consists of 20k examples, and evaluations are carried out on the full development and test sets. For mathematical reasoning tasks, we train models on DAPO-Math-17k-Processed (Yu et al., 2025), comprising 17k examples, and conduct evaluation on five test sets: DAPO-AIME-2024 (AI-MO, 2024; Bytedance & Tsinghua-SIA, 2025), AIME-2025 (Lin, 2025), MATH500 (Lightman et al., 2024; HuggingFaceH4, 2023), AMC23 (AI-MO, 2024), and Olympiad (He et al., 2024).
Baselines and Training Details. We use Qwen2.5-7B-Instruct and Qwen2.5-14B-Instruct as base models and Qwen2.5-72B-Instruct as the large LLM to generate GT Traj. The reference reward $R^{\text{ref}}$ is updated every 500 steps. For training, we set the learning rate to 1e-6, maximum response length to 8192, sampling temperature to 1.0 and top-p to 1.0. For inference, we set the sampling temperature to 0.6 and top-p to 0.95. For the multi-hop QA tasks, we benchmark our method against not only state-of-the-art LLMs such as DeepSeek-R1-0528, GPT-4.1-0414, O4-mini-0416, and Gemini-2.5-pro-0325, but also prominent RL-based agentic search models (Jin et al., 2025; Song et al., 2025). We adopt the ReSearch (Chen et al., 2025) framework, with pre-samples $M=5$, rollout $N=5$, train batch size of 8, and 1,000 training steps. For DynaSearcher(Hao et al., 2025), we remove the “kg_filter” during inference. For mathematical reasoning tasks, we primarily adopt GRPO (Shao et al., 2024), DAPO (Yu et al., 2025), and GSPO (Zheng et al., 2025) as baselines. We use the verl (Sheng et al., 2025) framework with pre-samples $M=16$, rollout $N=16$, train batch size of 32, and 1,000 training steps. For DAPO, we set the clipping parameter $\epsilon_{\text{low}}=0.2$ and $\epsilon_{\text{high}}=0.28$. For GRPO, we set the “loss_agg_mode” to “seq-mean-token-mean”, which is aligned with the original paper. For GSPO, the clipping parameter $\epsilon$ is set to 0.0003. All experiments are conducted on a server equipped with an Intel(R) Xeon(R) Platinum 8369B CPU and $8\timesNVIDIA$ A100-SXM4-80GB GPUs. More details can be found in Appendix A.1.
@@ 6 Experiments
In this section, we conduct a series of experiments to comprehensively evaluate PVPO. First, we test our method on multi-hop QA to validate its effectiveness in the agent domain. Next, we perform ablation studies to examine the contributions of the core modules of PVPO. We further apply PVPO to mathematical reasoning tasks to verify its generalizability and also evaluate its compatibility with other advanced RL algorithms. In addition, we analyze the training efficiency and convergence properties of PVPO. Finally, we present a case study to investigate the efficiency and robustness of PVPO under low sampling budget.
Table 1: Performance comparisons between PVPO and the baselines on multi-step retrieval datasets. The best and second best results are bold and underlined, respectively.
Method Musique 2Wiki HotpotQA Bamboogle Average
@@
Qwen2.5-7B-Instruct 5.1 13.5 27.9 29.3 22.4 31.0 12.8 17.1 17.1 22.7
DeepSeek-R1 32.0 40.7 57.5 59.4 43.0 58.3 66.4 76.6 49.7 58.8
O4-mini 38.0 44.1 61.5 67.4 49.5 67.4 74.4 84.2 55.9 65.8
GPT-4.1-global 31.0 40.9 58.0 58.5 44.5 57.7 51.2 61.6 46.2 54.7
Gemini-2.5-pro 42.5 50.8 70.0 71.2 53.0 71.1 75.2 84.5 60.2 69.4
@@
Search-R1-v0.3 24.7 34.6 58.7 61.1 53.6 66.9 48.0 54.5 46.3 54.4
R1-Searcher 24.7 34.2 67.8 68.2 59.7 71.5 46.4 52.0 50.5 56.5
GRPO-ReSearch 33.4 46.7 60.8 67.0 54.5 63.7 45.6 54.4 48.6 58.0
GRPO-DynaSearcher 38.9 52.0 74.3 76.8 62.7 68.3 51.2 58.7 56.8 64.0
PVPO-ReSearch 36.5 51.4 70.1 72.4 65.5 72.3 45.6 54.3 54.4 62.6
PVPO-DynaSearcher 46.9 59.4 77.7 80.6 69.0 78.4 50.4 59.7 61.0 69.6
6.1 Main Results
We evaluate PVPO against both zero-shot leading LLMs and trained RL-based search methods, with results in Table 1 underscoring its effectiveness. Specifically, applying PVPO substantially improves the base frameworks, boosting ReSearch’s Avg Acc/LasJ scores by 5.8/4.6 points and DynaSearcher’s by 4.2/5.6 points. Notably, our PVPO-DynaSearcher model significantly outperforms all RL-trained baselines (e.g., surpassing GRPO by over 5 points on average). It also marginally exceeding the strongest proprietary LLM, Gemini-2.5-Pro, while establishing a considerable lead over other models like O4-mini, GPT-4.1, and DeepSeek-R1. On the Bamboogle dataset, SOTA LLMs significantly outperform 7B-trained models largely due to the outdated 2018 Wikipedia corpus used in our experiments (see Appendix A.1 and Figure 5). Overall, these results demonstrate that PVPO consistently achieves state-of-the-art performance across agentic search methods.
6.2 Ablation Study
We conduct an ablation study to isolate the contribution of each component in PVPO, as shown in Table 2. Starting from the GRPO-DynaSearcher baseline (56.8 Avg Acc / 64.0 LasJ), the integration of Static V Estimation first raises the scores to 58.3/66.7. Subsequently adding Group Sampling further boosts the performance to 61.0/69.6, which represents our full PVPO model and outperforms all baselines. This incremental improvement validates the effectiveness of each proposed component.
Table 2: Ablation study of PVPO on multi-step retrieval datasets. Starting from DynaSearcher on Qwen2.5-7B-Instruct, we incrementally add Static V Estimation and Group Sampling.
Method Average
Acc LasJ
Qwen2.5-7B-Instruct 31.1 39.7
GRPO-DynaSearcher 56.8 64.0
+ Static V Estimation 58.3 66.7
+ Group Sampling 61.0 69.6
Figure 2: Training efficiency of PVPO on mathematical reasoning datasets.
Table 3: Performance comparison of PVPO and baseline methods on mathematical reasoning datasets using different model scales. “w/” means trained with.
Method MATH500 AMC23 Olympiad AIME-2024 AIME-2025 Avg Acc
Qwen2.5-7B-Instruct 75.68 42.92 38.94 12.10 6.67 35.26
w/ GRPO 78.60 49.10 42.14 13.86 10.10 38.76
w/ DAPO 78.58 51.38 43.36 14.96 11.30 39.92
w/ GSPO 78.66 50.12 43.60 15.02 12.70 40.02
w/ PVPO 80.30 52.02 44.62 14.86 14.70 41.30
Qwen2.5-14B-Instruct 79.68 51.52 44.00 14.82 12.29 40.46
w/ GRPO 82.12 53.50 47.42 16.14 15.86 43.01
w/ DAPO 82.50 56.44 49.34 18.04 15.66 44.40
w/ GSPO 83.56 56.02 49.28 18.18 16.20 44.65
w/ PVPO 83.64 56.78 50.72 19.24 17.74 45.62
@@ 6.3 Generalization Evaluation
To evaluate the transferability of PVPO, we apply it to mathematical reasoning tasks spanning a range of difficulties, from basic arithmetic to olympiad-level problems. We compare PVPO with GRPO, DAPO, and GSPO across several benchmark datasets. As shown in Table 3, PVPO consistently outperforms all baselines on both the 7B and 14B model scales. We further combine PVPO with the core modules of advanced RL methods, such as the sequence-level importance ratio from GSPO and the KL removal strategy from DAPO, achieving additional performance improvements when integrated with these state-of-the-art algorithms. Since these integrated modules are not the main focus of PVPO, we provide the detailed results and metrics for these extensions in Appendix A.3 and Table 4. Furthermore, PVPO exhibits robust cross-domain generalization and enhanced scalability.
@@ 6.4 Training Efficiency Analysis
[…]
Figure 3: Training stability of PVPO on multi-step retrieval datasets.
6.5 Stability Evaluation
We track PVPO training metrics to show its stability. Figure 3 (a) shows that PVPO achieves a much higher average reward than GRPO. With a similar KL divergence in Figure 3 (b), this improvement comes not from more aggressive updates, but from better gradient direction estimates. As shown in Figure 3 (c), PVPO has lower advantage variance, leading to more reliable and consistent update directions. PVPO also maintains exploration without losing stability. Figure 3 (d) shows that it keeps higher policy entropy under a similar KL constraint, which helps avoid premature convergence to a local optimum. Overall, PVPO addresses key problems in RL by supporting high exploration, low variance, and high rewards, thereby achieving more stable training than existing methods.
6.6 Case Study: Low Sampling Budget
To further examine PVPO’s performance under resource constraints, we conduct a case study on low sampling budget. We reduce the number of rollouts from 5 (used in the main experiments) to 2. For comparison, we report GRPO’s performance with a full budget. Figure 4 (a) shows that PVPO with a low budget remains close to the fully budgeted GRPO. We calculate computational cost by multiplying the number of rollouts with the average number of tool calls in trajectories. As shown in Figure 4 (b), PVPO’s average cost is only 4.3, which is much lower than GRPO’s 11.7. PVPO achieves 97% of GRPO’s performance (55.0% vs 56.8%) while using less than 40% of the computational cost. This strong sample efficiency comes from the high-quality, low-variance training signals provided by Static V Estimate. The model can update its policy efficiently using fewer rollouts.
Figure 4: Low sampling budget of PVPO on multi-step retrieval datasets. The NN denotes the number of trajectories in each single rollout. NN=5 is the full budget and NN=2 is the low budget.
7 Conclusions
In this paper, we propose PVPO, an efficient critic-free reinforcement learning algorithm designed to optimize policy learning for complex tasks. By introducing a Static V Estimate as an external advantage reference and integrating it with group sampling for effective data filtering, PVPO addresses the limitations of extensive sampling and biased intra-group comparisons inherent in prior methods. Our approach yields stable, low-variance training signals, accelerates convergence, and significantly reduces computational costs. Extensive experiments across nine diverse benchmarks in multi-hop QA and mathematical reasoning demonstrate that PVPO achieves state-of-the-art performance and strong generalization, even with small-scale models and limited resources. PVPO introduces substantial improvements in reasoning and tool use, supports scalable training, and ensures consistent performance, thereby demonstrating strong potential for widespread real-world application.
Appendix A Appendix
A.1 Implementation Details
Retriever and Corpus. For the multi-hop QA task, we employ multilingual-e5-base as the retriever model and use the December 2018 Wikipedia dump as the primary retrieval corpus, which contains over 21 million passages. To improve retrieval efficiency, we construct the final corpus by combining supporting document passages from three multi-hop datasets (i.e., Musique, 2Wiki, and HotpotQA) with one million randomly sampled documents from the Wikipedia dump. Notably, Bamboogle only provides questions and answers without ground truth passages, so it cannot be incorporated into the retrieval corpus. This may contribute to the lower scores on Bamboogle for most methods, as shown in Table 1.Passage retrieval is implemented using FAISS [^1] https://pypi.org/project/faiss-gpu/, and for each query, the top 5 passages are retrieved during both training and testing. For the KG (Knowledge Graph) data used in PVPO-DynaSearcher, we follow the approach and dataset provided by Wang et al. (2021), which is aligned with Hao et al. (2025).
Prompts and Code. We implement PVPO-ReSearch and PVPO-DynaSearcher based on the ReSearch framework [^2] https://github.com/Agent-RL/ReCall/tree/re-search. The system prompts for ReSearch and DynaSearcher are set following their respective original papers, detailed prompt templates are shown in Figure 6 and  7. For prompt-based SOTA LLMs, we first retrieve 5 passages from the corpus for each question, and then organize these passages using the template shown in Figure 5 as the prompt for answer generation. For mathematical reasoning tasks, we use verl version 0.3.1.dev0. Since the ReSearch codebase is also developed on top of the verl framework, we provide the core implementation of our PVPO method based on verl in code Listing 1.
1# verl/trainer/ppo/ray_trainer.py
2def compute_advantage(...):
3 if adv_estimator == AdvantageEstimator.PVPO:
4 # compute pvpo advantages
5 advantages, returns = core_algos.compute_pvpo_outcome_advantage(
6 token_level_rewards=data.batch["token_level_rewards"],
7 token_level_values=data.non_tensor_batch["static_value"],
8 response_mask=data.batch["response_mask"],
9 )
10 data.batch["advantages"] = advantages
11 data.batch["returns"] = returns
12 ...
13# verl/trainer/ppo/core_algos.py
14def compute_pvpo_outcome_advantage(
15 token_level_rewards: torch.Tensor,
16 token_level_values: torch.Tensor,
17 response_mask: torch.Tensor,
18):
19 scores = token_level_rewards.sum(dim=-1)
20 values = torch.tensor(token_level_values.astype(np.float32), device=scores.device, dtype=scores.dtype)
21
22 with torch.no_grad():
23 for i in range(scores.shape[0]):
24 scores[i] = (scores[i] - values[i])
25 scores = scores.unsqueeze(-1) * response_mask
26 return scores
Listing 1: PyTorch-style pseudocode for PVPO
[…]
Figure 5: Prompt for zero-shot LLM RAG.
[…]
Figure 6: System prompt for ReSearch.
[…]
Figure 7: System prompt for DynaSearcher.
You will be provided with three pieces of content: the questioner’s question, the user’s response, and the reference answer list. Your task is to score the accuracy of the user’s response based on the criteria outlined below. Please ensure that you carefully read and understand these instructions. Evaluation Criteria: 1. The pred answer doesn’t need to be exactly the same as any of the ground truth answers, but should be semantically same for the question. 2. Each item in the ground truth answer list can be viewed as a ground truth answer for the question, and the pred answer should be semantically same to at least one of them. 3. The user’s response may be longer and more detailed; as long as it is logically correct, contains the correct answer, it should be scored appropriately. Evaluation Steps: 1. Carefully read the questioner’s question and understand its key points. 2. Carefully read the reference answer and understand the key points relevant to the question. 3. Based on the evaluation criteria, assign a score in the range of 0 to 5, where 0 indicates that the user’s response does not include any of the key points from the reference answer and completely fails to answer the questioner’s question; 5 indicates that the user’s response includes all the key points from the reference answer and fully and correctly answers the questioner’s question. Questioner’s question: {question} Reference answer: {answer} User’s response: {response} Evaluation result (output only the score between 0 and 5):
Figure 8: Prompt for LLM-as-Judge score.
A.2 Group Sampling Analysis
[…]